OpenStudy (anonymous):

How do i do this? Graph -x+2y+2z=6

OpenStudy (lukecrayonz):

Have you ever graphed 3D?

OpenStudy (snowfire):

Draw 3 axes instead of 2 like you're used to, and label them x, y, and z. Then you need to use substitution to get your values for each variable.

OpenStudy (snowfire):

So first I would solve for x in terms of y and z by isolating it. x=2y+2z-6

OpenStudy (snowfire):

Then do the same and solve for y next in terms of z. Once you have that answer, solve for z now that you have all z's and then proceed to plug in your number value for z into the previous equations for x and y.

OpenStudy (ranga):

x = 0 for all points on the yz plane. y = 0 for all points on the xz plane. z = 0 for all points on the xy plane. So set x = 0 in you equation and you will get the equation of a straight line in the yz plane. Draw that line. Do the same on the other two planes. The graph of -x+2y+2z=6 will be a plane containing all those three lines drawn earlier.

OpenStudy (snowfire):

Actually I think my method is broken here, just ignore what I said.

OpenStudy (ranga):

-x+2y+2z=6 set x = 0. 2y+2z = 6, y + z = 3. Draw this line on the yz plane set y = 0. -x + 2z = 6. Draw this line on the xz plane set z = 0. -x + 2y = 6. Draw this line on the xz plane

OpenStudy (ranga):

Just thought of something: You can simplify the graphing even further by setting two of the coordinates at a time to zero. -x+2y+2z=6 Set x = 0, y = 0: 2z = 6 ; z = 3 This is where the plane cuts the z axis. Set x = 0, z = 0: 2y = 6 ; y = 3 This is where the plane cuts the y axis. Set y = 0, z = 0: -x = 6 ; x = -6 This is where the plane cuts the y axis. Knowing three points, is enough to determine the plane. So draw a plane passing through those three points.  The sketch does not depict the entire plane but just shows where the plane intersects the x,y,z axes and the xy, yz, xz planes.